# Magnetosphere Evolution and Precursor-Driven Electromagnetic Signals in Merging Binary Neutron Stars

### v1

@@ 4.4 Post-Merger Signals and Afterglow Echos
Because the EM energy escaping a BNS system is strongly anisotropic, most observers will not be situated where the bulk of the outflow is directed. However, even if an observer’s line of sight does not coincide with the initial direction of the outflow, interactions with the surrounding medium may lead to secondary or echo signals. In particular, magnetically dominated precursor outflows could generate a pulsar-wind-like shock (e.g., N. Sridhar et al., 2021) before the actual merger, modifying the surrounding circumburst medium well in advance of the final coalescence. This interaction may later contribute to delayed EM signals that are distinct from standard post-merger afterglow components (e.g., R. Sari, 1998; J. E. Rhoads, 1999; J. Granot et al., 2002).
@@
Unlike standard GRB afterglow echoes, where a delayed signal arises due to the expansion and deceleration of a jet-aligned outflow, pre-merger EM outflows may be emitted in directions distinct from both the observer’s line of sight and the GRB jet. If these precursor-driven outflows later interact with a dense region of circumburst material, they can decelerate, expand, and reprocess their energy, leading to secondary radiation that eventually reaches the observer with a delay relative to the GW signal. The total observed delay of such an echo consists of two components: a geometric delay due to the additional distance traveled before encountering the medium and a dynamical delay associated with the deceleration and expansion of the outflow.
@@
Such precursor-driven echoes may contribute to late-time re-brightening in short GRBs. For instance, GW170817 exhibited prolonged X-ray afterglow re-brightening up to 1000 days post-merger (J. J. Ruan et al., 2018; E. Troja et al., 2020), and some short GRBs have shown late-time radio activity of unknown origin (e.g., G. Schroeder et al., 2024). While these events are often attributed to prolonged central engine activity or structured jets, pre-merger EM outflows could present an alternative or complementary explanation. Notably, precursor-driven echoes could also appear as orphan afterglows associated with GW events at unusually large inclination angles.
Yet several uncertainties persist. The presence of a sufficiently strong pre-merger outflow requires at least one highly magnetized NS. The environment at off-axis angles must allow efficient shock formation, and the resulting emission must be strong enough to be detectable. Hence, while pre-merger echoes offer an intriguing explanation for some late-time GRB re-brightenings, their actual contribution and detectability remain to be tested by future deeper multi-wavelength observations, particularly in the radio and X-ray bands.
@@ 5 Summary and Future Directions
In this study, we performed FF simulations of BNS magnetospheres, exploring their evolution, EM outflows, and potential observational signatures. We analyzed the structure of the Poynting flux, high-energy emission distributions, and the development of EM torques, which may have subtle but cumulative effects on the orbital motion. While our results provide insight into the fundamental mechanisms shaping magnetospheric interactions, several aspects remain to be explored.
@@
In particular, our analysis neglects tidal distortions and magnetosphere-induced forces on the NS surfaces, which could contribute to additional dynamics, such as crustal deformations or resonant crust-shattering events. Although these effects are unlikely to alter the qualitative structures derived here, they may influence the fine details of precursor EM signals. Future studies should incorporate these effects self-consistently, alongside kinetic plasma simulations and more advanced MHD treatments, to refine the modeling of pair production, particle acceleration, and radiation transport. Additionally, incorporating full general relativistic effects will be necessary to capture the intricate coupling between magnetospheric evolution and the strong-gravity environment near the merger. These improvements will be essential for more accurate predictions of precursor signals and their detectability in upcoming multi-messenger observations.
@@
The material is based upon work supported by NASA under award numbers 80GSFC21M0002 and 80GSFC24M0006, 80NSSC21K1999, 22-ADAP22-0142, 22-TCAN22-0027, 21-ATP21-0116, and 22-FERMI22-0035. We acknowledge insightful discussions with Ioannis Contopoulos, Paz Beniamini, Lauren Rhodes, and Vikram Ravi. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. This work has made use of the NASA Astrophysics Data System. For the analysis and visualization of the simulation results Mathematica (W. R. Inc., 2024), Matplotlib (J. D. Hunter, 2007), Plotly (P. T. Inc., 2015), PyVista (C. B. Sullivan & A. Kaszynski, 2019), CMasher(E. van der Velden, 2020), Cartopy (Met Office, 2010 - 2015), NumPy (C. R. Harris et al., 2020), Numba (S. K. Lam et al., 2015) and Scipy (P. Virtanen et al., 2020) are used.

### v2

@@ 4.4 Postmerger Signals and Afterglow Echos
Because the EM energy escaping a BNS system is strongly anisotropic, most observers will not be situated where the bulk of the outflow is directed. However, even if an observer’s line of sight does not coincide with the initial direction of the outflow, interactions with the surrounding medium may lead to secondary or echo signals. In particular, magnetically dominated precursor outflows could generate a pulsar-wind-like shock (e.g., N. Sridhar et al., 2021) before the actual merger, modifying the surrounding circumburst medium well in advance of the final coalescence. This interaction may later contribute to delayed EM signals that are distinct from standard postmerger afterglow components (e.g., R. Sari, 1998; J. E. Rhoads, 1999; J. Granot et al., 2002).
@@
Unlike standard GRB afterglow echoes, where a delayed signal arises due to the expansion and deceleration of a jet-aligned outflow, premerger EM outflows may be emitted in directions distinct from both the observer’s line of sight and the GRB jet. If these precursor-driven outflows later interact with a dense region of circumburst material, they can decelerate, expand, and reprocess their energy, leading to secondary radiation that eventually reaches the observer with a delay relative to the GW signal. The total observed delay of such an echo consists of two components: a geometric delay due to the additional distance traveled before encountering the medium and a dynamical delay associated with the deceleration and expansion of the outflow.
@@
Such precursor-driven echoes may contribute to late-time rebrightening in short GRBs. For instance, GW170817 exhibited prolonged X-ray afterglow rebrightening up to 1000 day postmerger (J. J. Ruan et al., 2018; E. Troja et al., 2020), and some short GRBs have shown late-time radio activity of unknown origin (e.g., G. Schroeder et al., 2024). While these events are often attributed to prolonged central engine activity or structured jets, pre-merger EM outflows could present an alternative or complementary explanation. Notably, precursor-driven echoes could also appear as orphan afterglows associated with GW events at unusually large inclination angles.
Yet, several uncertainties persist. The presence of a sufficiently strong premerger outflow requires at least one highly magnetized NS. The environment at off-axis angles must allow efficient shock formation, and the resulting emission must be strong enough to be detectable. Hence, while premerger echoes offer an intriguing explanation for some late-time GRB rebrightenings, their actual contribution and detectability remain to be tested by future deeper multiwavelength observations, particularly in the radio and X-ray bands.
@@ 5 Summary and Future Directions
In this study, we performed FF simulations of BNS magnetospheres, exploring their evolution, EM outflows, and potential observational signatures. We analyzed the structure of the Poynting flux, high-energy emission distributions, and the development of EM torques, which may have subtle but cumulative effects on the orbital motion. In particular, by simulating a wide range of magnetic configurations, we confirmed several robust features of BNS magnetospheric interactions while also uncovering novel trends. We systematically quantified the strong anisotropy of the Poynting flux distributions and established, for the first time, a configuration-dependent power-law relationship between the Poynting flux luminosity and the orbital frequency. The power-law index was found to vary significantly with the relative alignment and strength of the stellar magnetic moments. Furthermore, we introduced a new method to construct high-energy emission skymaps and light curves for BNS systems by generalizing techniques originally developed for isolated pulsars. Finally, our analysis of the photon spectra revealed that, although extremely high-energy photons (up to TeV-PeV) can be produced in the radiation reaction limit, magnetic pair production severely attenuates them. As a result, the escaping emission is constrained to the MeV range, identifying this as a promising observational window for precursor high-energy signals. While our results provide insight into the fundamental mechanisms shaping magnetospheric interactions, several aspects remain to be explored.
@@
In particular, our analysis neglects tidal distortions and magnetosphere-induced forces on the NS surfaces, which could contribute to additional dynamics, such as crustal deformations or resonant crust-shattering events. Although these effects are unlikely to alter the qualitative structures derived here, they may influence the fine details of precursor EM signals. Future studies should incorporate these effects self-consistently, alongside kinetic plasma simulations and more advanced MHD treatments, to refine the modeling of pair production, particle acceleration, and radiation transport. Additionally, incorporating full general relativistic effects will be necessary to capture the intricate coupling between magnetospheric evolution and the strong-gravity environment near the merger. These improvements will be essential for more accurate predictions of precursor signals and their detectability in upcoming multimessenger observations.
@@
We would like to thank the anonymous referee for useful comments that helped to improve the manuscript. We also acknowledge insightful discussions with Ioannis Contopoulos, Paz Beniamini, Lauren Rhodes, and Vikram Ravi. The material is based upon work supported by NASA under award numbers 80GSFC21M0002 and 80GSFC24M0006, 80NSSC21K1999, 22-ADAP22-0142, 22-TCAN22-0027, 21-ATP21-0116, and 22-FERMI22-0035. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. This work has made use of the NASA Astrophysics Data System. For the analysis and visualization Mathematica (W. R. Inc., 2024), Matplotlib (J. D. Hunter, 2007), Plotly (P. T. Inc., 2015), PyVista (C. B. Sullivan & A. Kaszynski, 2019), CMasher(E. van der Velden, 2020), Cartopy (Met Office, 2010 - 2015), NumPy (C. R. Harris et al., 2020), Numba (S. K. Lam et al., 2015) and Scipy (P. Virtanen et al., 2020) are used.
Appendix A Resolution Tests and Robustness of Results
Figure 29: The Poynting flux evolution for the cases with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\Delta\Phi=0^{\circ}$, and $\lambda=1$ (
$\uparrow$
$\uparrow$
) in red, and $\lambda=-1$ (
$\uparrow$
$\downarrow$
) is in black with the regular (solid line) and double (dashed line) resolution.
Figure 30: Similar to Figure 8 the structure of the magnetosphere, at $\tau\simeq-7ms,$ for the cases with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\Delta\Phi=0^{\circ}$, and $\lambda=1$ (
$\uparrow$
$\uparrow$
) in top row while $\lambda=-1$ (
$\uparrow$
$\downarrow$
) in bottom row. Left are the simulations with the resolution used in the main text while right are the higher-resolution simulations.
Figure 31: Similar to Figure 26. Snapshots of the evolution of high-energy emission skymaps for the case with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\Delta\Phi=0^{\circ}$, and $\lambda=-1$ (
$\uparrow$
$\downarrow$
) in higher resolution.
To assess the robustness of our results, higher resolution were conducted for selected cases. For these simulations, a double spatial and time resolution was employed relative to the baseline simulations. Specifically, the grid size was reduced to $0.05r_{\star}$ - half compared to the one used for the main results - and the time step to $dt=0.02r_{\star}/c$ in order to maintain the same Courant number.
In Figure 29, the Poynting flux evolution is shown for the cases with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\Delta\Phi=0^{\circ}$, and $\lambda=1$ ( $\uparrow$ $\uparrow$ ) in red while $\lambda=-1$ ( $\uparrow$ $\downarrow$ ) is in black. The solid lines correspond to the simulations with the baseline resolution used in the rest of this study, while the dashed lines correspond to simulations with double the resolution.
In both cases, the Poynting flux evolution obtained from the higher-resolution simulations is in great agreement with that of the baseline simulations presented in the main text, indicating that our main results are not sensitive to the resolution. As expected from such flux curves, the values of the power-law index of the higher-resolution runs are indistinguishable in both resolutions, $3.3$ and $2.9$ for the $\uparrow$ $\uparrow$ and $\uparrow$ $\downarrow$ case, respectively.
To further verify the fidelity of our adopted resolution in the main text, we present in Figure 30, the magnetospheric structure at $\tau\simeq-7ms$ for both configurations across both resolutions. In addition, Figure 31 compares the evolution of the high-energy emission skymaps from both resolutions of the ( $\uparrow$ $\uparrow$ ) case. Lastly, Figure 32 displays the evolution of the EM torque for both resolutions.
Across all diagnostics, the global magnetospheric structures exhibit only minor differences. The torque evolution is nearly identical, further supporting the convergence of our results. The high-energy emission skymaps, while exhibiting some small discrepancies, particularly in finer details, remain consistent in their overall morphology and dominant features. These differences are understandable, as the skymap construction depends on localized regions of strong accelerating electric fields, which are more sensitive to resolution-dependent magnetic topology variations.
Overall, the higher-resolution simulations validate the accuracy and robustness of our primary results. The modest differences observed suggest that the resolution used in the main analysis is sufficient to capture the essential features without incurring significantly higher computational cost.
H Refer to caption
Figure 32: The evolution of the torque for the case with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\Delta\Phi=0^{\circ}$, and $\lambda=-1$ (
$\uparrow$
$\downarrow$
) in the regular (solid line) and higher resolution (dashed line).